OpenStudy (anonymous):

you need to know the number of different arrangements possible for five distinct letters. You decide to use the permutations rule,, but your friend tells you to use 5. who is correct and why????? please help me ive been on this question for 30 mins

OpenStudy (kropot72):

There are 5 choices for the first letter. Having chosen a first letter there are 4 choices for the second letter. Therefore there are 5 * 4 = 20 choices for the first two letters. Continuing this reasoning gives the following number of arrangements for the five letters taken five at a time: \[5\times4\times3\times2\times1=5!\] If we use the formula for permutations we get \[5P5=\frac{5!}{(5-5)!}=5!\]

OpenStudy (kropot72):

Remember, in statistics, 0! = 1

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)